Present Value Years Interest Rat Future Value Equation
13 9% $16,832
4 7% $48,318
29 13% $886,073
40 21% $550,164

For each of the following, compute the present value. (round answers to two decimal places)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.

16832=P*(1.09)^13

P=16832/(1.09)^13

=$5490.24(Approx).

2.

48318=P*(1.07)^4

P=48318/(1.07)^4

=$36861.57(Approx).

3.

886,073=P*(1.13)^29

P=886,073/(1.13)^29

=$25597.33(Approx).

4.

550164=P*(1.21)^40

P=550164/(1.21)^40

=$268.58(Approx).